please help me solve this equation -(2x+9)-(6x-2)=7-(2x-10)
Answer by fractalier(6550)

Okay, from
-(2x + 9) - (6x - 2) = 7 - (2x - 10)
let us remove parentheses by distributing minus signs and collect like terms...
-2x - 9 - 6x + 2 = 7 - 2x + 10
-8x - 7 = -2x + 17
now add 7 to both sides and then add 2x to both sides
-8x = -2x + 24
-6x = 24
now divide by -6
x = -4